OpenStudy (anonymous):

the sum of the digits of a two- digit number is 7. If the digits are reversed the number is increased by 27. What is the number?

OpenStudy (anonymous):

let the number be xy. Then x+y=7 and 10y+x=10x+y+27. So just solve this system of two equations

OpenStudy (anonymous):

It makes sense now. I just couldn't get it set up. Thank you.

OpenStudy (jiteshmeghwal9):

Let first digit be x.Then, second digit of the number = (7-x) Original number=10*(7-x)+x=70-10x+x=70-9x reversed number=10*x+(7-x)=10x+7-x=9x+7 A/q, => (70-9x)+27=9x+7 solve for x now :)

OpenStudy (anonymous):

sum of digits is x+y=7 -equation (1) the number will be 10x+y if we chnge the place of digits the nmber increases by 27 so 10x+y=10y+x-27 => 9x-9y=-27 => x-y=-3 -equation(2) sum up equation 1 nd 2 => 2x=4 x=2 putting value of x in eq(1) y=5 so the number will be 10(x)+y =20+5 =25 .............. now its digit sum is 2+5=7 nd change its digits place it will be 52 which is 27 greater than 25...

OpenStudy (jiteshmeghwal9):

@khurramshahzad is right also. only difference is that he did the solution using two variables & I did with one variable :)

OpenStudy (anonymous):

yup

OpenStudy (jiteshmeghwal9):

:)

OpenStudy (amoodarya):

 your problem has 6 answer , I list in a draw